OpenStudy (anonymous):

Let f(x) = 2x+3 and g(x) = x2 Part 1 Find f(g(x)) Part 2 Find g(f(x)) Part 3 Use complete sentences to explain any difference in part 1 and part 2.

OpenStudy (anonymous):

1. 2x^2 + 3 2.4x^2+ 9 +12x

OpenStudy (akshay_budhkar):

2x^2+3 for first 4x^2+9+12x

OpenStudy (anonymous):

\[f\circ(g(x))=f(g(x))=f(x^2)=2x^2+3\] \[g(f(x))=g(2x+3)=(2x+3)^2\]

OpenStudy (anonymous):

Are there any differences in part one and part two? If so, please explain. This is extremely important.

OpenStudy (anonymous):

much different. first one says 1) first square, 2) then multiply by 2 and add 3 second one says 1) first multiply by 2 and add 3 2) then square the result

OpenStudy (anonymous):

so for example if x was 10, first one would say \[10^2=100,\text{then }2\times 100+3=203\] second one says \[2\times 10+3=23,\text{ then } 23^2=529\] they are totally different

OpenStudy (anonymous):

If f(x) = x + 2, find f(3).

OpenStudy (anonymous):

????